Suppose that a simple random sample is taken from a normal population having a standard deviation of 11 for the purpose of obtaining a 95​% confidence interval for the mean of the population. a. If the sample size is 16​, obtain the margin of error. b. Repeat part​ (a) for a sample size of 81.

a. The margin of error for a sample size of 16 is ??? ​(Round to two decimal places as​ needed.)

b. The margin of error for a sample size of 81 is ???
Answer #1

SolutionA:

z alpha/2 for 95%=1.96

n=16

population standard deviation =sigma=11

margin of error=zcrit*sigma/sqr(n)

=1.96*11/sqrt(16)

=1.96*11/4

margin of error =5.39

ANSWER:5.39

Solutionb:

z alpha/2 for 95%=1.96

n=81

population standard deviation =sigma=11

margin of error=zcrit*sigma/sqr(n)

=1.96*11/sqrt(81)

=2.396

margin of error=2.4

ANSWER:2.4